OpenStudy (anonymous):

Please help me with the following: Solve: (5sec3x-2) (sec3x+1) =0 within [0,2Pi)

OpenStudy (tkhunny):

Solve: 5sec(3x) - 2 = 0

OpenStudy (anonymous):

?? confused

OpenStudy (tkhunny):

Remember when yousolved quadratic equations? x^2 - x - 2 = 0 Factor, if possible (x-2)(x+1) = 0 If we are ringing any bells, here, we need to ask, "What did you do next to solve this quadratic equation?"

OpenStudy (anonymous):

set each side = to 0 and solve